Craig, Inc., has provided the following information for one of its products for each hour of production: Actual velocity: 100 units (per hour) Move time: 20 minutes Inspection time: 14 minutes Rework time: 13 minutes Required:

1. Calculate MCE. If required, round your answer to two decimal places.

2. What is the theoretical cycle time? Calculate MCE using actual and theoretical cycle times. If required, round your answers to two decimal places. Theoretical cycle time minutes Actual cycle time minutes MCE

3. What if waste is reduced by 30 percent? New waste = minutes What is the new MCE? If required, round your answer to two decimal places. What is the new cycle time? minutes
Answer #1

Solution:

1. MCE=Value added time (or Process time) / Manufacturing cycle time

=60/(60+20+14+13)

= 100/107= 93.45%

2.Theoritical cycle time includes only time that devotes for value added activities in otherwords we can say it includes only process timings and excludes all handover or wait time.

Actual cycle time: 0.94 minutes/unit

Where as MCE at actual cycle time= 94/107 = 87.85%

Theoritical cycle time = 0.6 minutes/unit

Where as MCE at Theoritical cycle time= 60/107 = 56.07%

3. if waste time reduce by 30% then new cycle time = 1.03 minute/unit

and new MEC = 60/103.1

= 5819%